OpenStudy (anonymous):

A 50 kg block moving initially slideing on a level surface at 20 m/s is subject to a constant friction of 10N untill it comes to a stop. How far will it slide and how much energy is dissipated due to the friction

OpenStudy (matt101):

To know how far the block slides, we need to know it's acceleration in order to know when it stops. You can find the acceleration using Newton's Second Law: \[F=ma\]\[-10=50a\]\[a=-0.2\] So the acceleration of the block is -0.2 m/s^2. Remember, friction is acting in the direction opposite the motion (hence -10 in the equation), which means the acceleration is also negative (in fact, it's a deceleration). We have acceleration [a], initial speed [v(i)], final speed [v(f)] (it's 0 m/s because the block stops), and we want to find distance travelled. That means you can use this equation: \[v_f^2=v_i^2+2ad\]\[0=400-2(0.2)d\]\[d=1000\] That means it will take 1000 m for the block to stop moving. The energy dissipated by friction can be calculated either as the work done by friction on the block: \[W=Fd\]\[W=10(1000)\]\[W=10000\] Or as the kinetic energy of the block initially (because friction converts all this energy to heat): \[KE={1 \over 2}mv^2\]\[KE={1 \over 2}(50)(20)^2\]\[KE=10000\] Either way, you get 10 000 J as the energy dissipated by friction! If you have any questions about any of this please let me know!

OpenStudy (anonymous):

How did you get the 400

OpenStudy (matt101):

Initial velocity squared - initial velocity is 20 m/s in the question, 20 squared is 400.

OpenStudy (anonymous):

Oh sorry my bad

OpenStudy (matt101):

No problem!